Find the coordinates of the midpoint of a segment with the given endpoints:
P(-12,4)Q (-3,-6)

Answer by checkley71(8403)

X DISTANCE IS -3-(-12)=-3+12=9 & THE Y DISTANCE IS 4-(-6)=10
X MIDPOINT IS HALF OF THE DISTANCE OR 9/2=4.5 OR -12+4.5=7.5
Y MIDPOINT IS HALF OF THE DISTANCE OR 10/2=5 OR -6+5=-1
SO THE MIIDPOINT IS (7.5,-1)